OpenStudy (anonymous):

√2 cosxtanx−tanx=0 Solve for 0

OpenStudy (anonymous):

>:( OKAY. I need to factor.

Directrix (directrix):

√2 cosxtanx−tanx=0 tan(x) [ √2 cos(x) - 1 ] = 0 Zero Product Property Time

Directrix (directrix):

tan(x) [ √2 cos(x) - 1 ] = 0 tan(x) = 0 OR √2 cos(x) - 1 = 0 tan(x) = 0 or cos(x) = 1/√2 Choose x values from restricted domain 0<x<pi (radians)